The percent composition by mass of oxygen in BaSO4 is
  1. 233.4%
  2. 66.7%
  3. 27.4%
  4. 58.7%
  5. 13.7%

Grade:12

2 Answers

Arun
5 years ago
Molar mass of BaSO4 = 233.3896 g/mol
 
hence percentage of oxygen = 64/ 233.3896 = 0.274 * 100 = 27.41%
 
hopeit helps
 
Regards
Arun
 
Shivangi Khatter
5 years ago
dear student
the correct answer is(C)
Molecular weight of BaSO4 = 233.43
Percentage composition by mass of oxygen = 4 *16 *100/233.43 = 27.4%